OpenStudy (anonymous):

How to get the limit (without l'hopital) of (sin (x) - x)/(tan(x)-x) limit x approaches 0.

OpenStudy (anonymous):

\[\sin (x)=x-\frac{x^3}{6}+O\left(x^5\right)\\ \tan (x)=x+\frac{x^3}{3}+O\left(x^5\right)\]

OpenStudy (anonymous):

uh oh... i forget to try this with Taylor ,but i was hoping for something using only substitution or cancelling. any way thanks a lot man.

OpenStudy (anonymous):

The limit is \( - \frac 1 2\)

OpenStudy (anonymous):

Ah, my mistake!

OpenStudy (anonymous):

Because \[ \frac { \frac{-x^3}6} {\frac{x^3}{3}}=-\frac 1 2 \]

OpenStudy (anonymous):

Actually, using Series is like using derivatives and derivatives are used in L'Hospital's Rule

OpenStudy (anonymous):

Thus this is the only way to get it. Can't anyone find a simple way??

OpenStudy (xapproachesinfinity):

you can separate the fraction an manipulate it little bit to get the usual limit x/sinx tanx/x and 1/cosx and so on i guess the result is 1-1/2=-1/2

OpenStudy (anonymous):

Thanks a lot man.